OpenStudy (anonymous):

Medal&Fan The lengths of the sides of a triangle form three consecutive odd numbers. The perimeter is 21 centimeters. What is the length of the longest side? Select one: a. 7 cm b. 11 cm c. 9 cm d. 8 cm

hartnn (hartnn):

you know what consecutive odd numbers mean ?

OpenStudy (anonymous):

noooo

OpenStudy (whpalmer4):

11,13,15 would be consecutive odd numbers, for example

OpenStudy (tennistar):

Ok, odd numbers are numbers that are not even (5,7) Consecutive is in a row.

hartnn (hartnn):

consecutuve mean numbers following each other, like 2,3,4 , 100,101,102 but consecutive odd means none of them are divisible by 2 like 3,5,7,9...

OpenStudy (whpalmer4):

so if the first number is x, the second number is x+2 and the third number is x+4. The perimeter is the sum of those 3 quantities, and equals 21. Can you write and solve the equation that reflects that?

OpenStudy (anonymous):

no im still confused, ehh.

OpenStudy (whpalmer4):

well, do you agree that every other number is odd?

OpenStudy (anonymous):

umm sure.c:

OpenStudy (anonymous):

The answer is C. 9

OpenStudy (anonymous):

how

OpenStudy (whpalmer4):

Okay, so if we let x be the first number, then x+2 is the next number, right?

OpenStudy (anonymous):

yeahh

OpenStudy (whpalmer4):

and x+2 + 2 = x+4 is the 3rd number

OpenStudy (whpalmer4):

and if we add the three of them together, they equal the perimeter, which is 21: \[x+(x+2)+(x+4) = 21\]\[x+x+x + 2 + 4 = 21\]\[3x + 6 = 21\]\[3x+6-6 = 21-6\]\[3x=15\]\[\frac{3x}{3} = \frac{15}{3}\]\[x=5\]Now, \(x\) is the first of the 3 numbers, and we want the 3rd of the 3 numbers. As you may recall, the expression for the third number was \(x+4\), so the third number is \(5+4 = 9\) Or we could just count off 5, 7, 9...

OpenStudy (whpalmer4):

Checking our answer: 5+7+9 = 21

OpenStudy (anonymous):

dangg. i wish math came to me as easy as it comes to you. well thanks so muchh.:)

OpenStudy (whpalmer4):

I've had some practice ;-)

OpenStudy (anonymous):

oh I bet.c: